purpleowers66:

The coordinates of the endpoints of and are A(x1, y1), B(x2, y2), C(x3, y3), and D(x4, y4). Which condition proves that ?

MoonShadow:

Two lines are parallel if they have the same slope. Computing for the slope of line segment ab and line segment cd using m=(y2-y1)/x2-x1): m of ab=(2-1)/(2-1)=1         m of cd=(4-3)/(4-3)=1 Since line segment ab and line segment cd have the same slope which is 1, then they are said to be parallel. 

MoonShadow:

And your welcome